1. How much heat is required 200 mL of ice at - 22 degrees Celsius (at typical Freezer temperature).
  2. One mole of a monoatomic perfect gas initially at temperature T0 expands From volume V0 to 2V0 (a) at constant temperature, (b) at constant pressure. Calculate the work of expansion and the heat absorbed by the gas in each case.
  3. Determine the mass and specific volume of argon gas in a vessel at 150 kPa and 20 degrees Celsius. The vessel is spherical and has a radius of 5m.
  4. A container which has a volume 0.1 m3 is fitted with plunger enclosing 0.5 kg of steam at 0.4 MPa. Calculate the amount of heat transferred and the work done when the steam is heated to 300 degrees Celsius at constant pressure.
1
Expert's answer
2021-04-13T06:51:15-0400

Question 1 is unclear

Question 2

"\u0394U= \\frac{3}{2}\n\n\n\u200b\t\n P(V_{2}\n\n\u200b\t\n \u2212V _{1}\n\u200b\t\n )"


"= \n\\frac{3}{2}\n\u200b\t\n P(2V_{0}\n\u200b\t\n \u2212V _{0}\n\u200b\t\n )= \n\\frac{3}{2}\n\u200b\t\n PV _{0}\n\u200b\t\n = \n\\frac{3}{2}\n\u200b\t\n RT _{0}\u200b"


"W=P(V _{2}\n\u200b\t\n \u2212V _{0}\n\u200b\t\n )=PV_{0}\n\u200b\t\n =RT _{0}\n\u200b"

"\\therefore Q= \n\\frac{3}{2}\n\u200b\t\n RT \n\u200b\t\n +RT_{o}\n\u200b\t\n = \n\\frac{5}{2}\n\u200b\t\n RT _{0}\n\u200b"


Question 3

specific volume = "\frac{Volume}{mass}"


Specific volume ="\frac{R \cdot T } {P}"

where

R = ideal gas constant =  8.314 J mol−1 K−1

T= temperature = 20 + 273 = 293 K

P = pressure = 150kPa = 150000 Pa


Specific volume "=\\frac{8.314 \u00d7 293 } {150000} \n\n\u200b"


specific volume = 0.016 m3 kg-1


Mass = "\frac{volume}{s.volume}"


volume = "\frac{4}{3}πr^{3} = \frac{4}{3}π5^{3} = 523.8 m^{3}"


mass = "\frac{523.8}{0.016} = 32737.5 kg"


Question 4

Amount of heat = mcΔθ

m= mass = 0.5kg

c= specific heat capacity of steam = 2110 J/Kg•K

Δθ = 300 - 100 = 200 degree Celsius


Amount of heat = 0.5 × 2110 × 100 = 105500J


Work done = PΔV

P= pressure = 0.4 MPa = 400000Pa

ΔV = 0.15 - 0.10 = 0.05

"\frac{V_1}{T_1} = \frac{V_2}{T_2}"


"\frac{0.10}{373} = \frac{V_2}{573}"


V2 = 0.15

Work done = 400000 × 0.05 = 20000J